OpenStudy (anonymous):

Solve for x:4/5x+4/3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar. @sodukogirl13

OpenStudy (anonymous):

(4/5x)+(4/3)=2x -(4/5x) -(4/5x) (4/5x)=((1+1/5)x) (4/3)(5/6)=((1+1/5)x)(5/6) 1+1/9=x

OpenStudy (texaschic101):

just like any fractions, when adding them you need to have a common denominator. What is the common denominator for 5,3, and 1 ?

OpenStudy (anonymous):

so what sthe answer @sodukogirl13

OpenStudy (anonymous):

1+1/9

OpenStudy (texaschic101):

you want to learn how to do it or just get the answer ?

OpenStudy (anonymous):

i just want answers i know how to already

OpenStudy (texaschic101):

4/5x + 4/3 = 2x multiply the entire equation by the LCD which is 15 12x + 20 = 30x -- subtract 12x from both sides 20 = 30x - 12x 20 = 18x 20/18 = x 10/9 = x turned into a mixed number = 1 1/9

OpenStudy (texaschic101):

If you multiply the equation by the LCD then it will get rid of the fractions...or you can deal with the fractions and do it the way sodukogirl did it.